OpenStudy (anonymous):

MEDAL AND FAN. Give three functions in standard form. one function,f(x), with two real rational solutions. one function,g(h),with two real irrational solutions. one function,h(x), with two complex solutions.

OpenStudy (anonymous):

f(x) = (x-3)(x-2) g(x) = (x - sqrt(2)(x-sqrt(3)) h(x) = (x-2i)(x-3i)

OpenStudy (anonymous):

@TexasBoy123 what does or what is "sqrt" mean? /:

OpenStudy (anonymous):

Square root

OpenStudy (anonymous):

omg, i knew that! lol Thank youu! Your a life saver.

OpenStudy (anonymous):

Anytime!